# RobGC: Towards Robust Graph Condensation

### v1

@@ Abstract
Graph neural networks (GNNs) have attracted widespread attention for their impressive capability of graph representation learning. However, the increasing prevalence of large-scale graphs presents a significant challenge for GNN training due to their computational demands, limiting the applicability of GNNs in various scenarios. In response to this challenge, graph condensation (GC) is proposed as a promising acceleration solution, focusing on generating an informative compact graph that enables efficient training of GNNs while retaining performance. Despite the potential to accelerate GNN training, existing GC methods overlook the quality of large training graphs during both the training and inference stages. They indiscriminately emulate the training graph distributions, making the condensed graphs susceptible to noises within the training graph and significantly impeding the application of GC in intricate real-world scenarios. To address this issue, we propose robust graph condensation (RobGC), a plug-and-play approach for GC to extend the robustness and applicability of condensed graphs in noisy graph structure environments. Specifically, RobGC leverages the condensed graph as a feedback signal to guide the denoising process on the original training graph. A label propagation-based alternating optimization strategy is in place for the condensation and denoising processes, contributing to the mutual purification of the condensed graph and training graph. Additionally, as a GC method designed for inductive graph inference, RobGC facilitates test-time graph denoising by leveraging the noise-free condensed graph to calibrate the structure of the test graph. Extensive experiments show that RobGC is compatible with various GC methods, significantly boosting their robustness. It effectively addresses different types and levels of graph structural noise, underscoring its utility in enhancing the applicability of GC in real-world scenarios.
@@ I Introduction
Graph data [1, 2, 3] is extensively utilized across diverse domains, owing to its capability to represent complex structural relationships among various entities in the real world. To address the imperative for data management, graph neural networks (GNNs) have attracted significant attention for their exceptional representation capabilities. By incorporating node attributes and graph topology, they could extract latent information within graphs and have been utilized in a wide range of real-world applications, including chemical molecules [4], social networks [5], and recommender systems [6, 7]. However, the burgeoning volume of graph data within these real-world applications poses formidable challenges in training GNN models. GNNs commonly employ the message-passing mechanism that propagates node features across the graph structure. Despite its effectiveness, this method can lead to the neighbor explosion issue [8], which considerably escalates the computational requirements during the model training stage. This situation becomes even more challenging when training multiple GNN models, such as in the application of neural architecture search [9], continual learning [10, 11], and federated learning [12].
In response to the urgent demand for processing large-scale graph data, graph condensation (GC) [13, 14, 15] is proposed to substitute the large training graph with a compact (e.g., 1,000×1,000$\times$1 , 000 × smaller) yet informative graph to accelerate the GNN training. By retaining the crucial characteristics of the large training graph, GNN models trained on the condensed graph demonstrate comparable performance to those trained on the large training graph, while significantly reducing model training time. This not only facilitates large-scale graph management but also expands the application of GNNs in resource-constrained scenarios for various downstream tasks. In this paper, we specifically focus on the inductive inference task, which is practical and widely utilized in existing high throughput graph systems [16, 17, 18, 19]. In this setting, new and previously unseen nodes are continuously integrated into the existing training graph as depicted in Fig. 1. Subsequently, GNNs trained on condensed graphs are deployed on this expanded graph to generate embeddings for unseen nodes and make predictions [20, 21].
@@
Despite the potential of GC methods to expedite model training, their practical applicability remains sub-optimal in intricate real-world scenarios. Predominantly, existing GC methods presuppose the presence of clean training graph structures, contradicting the complicated nature of data processing in practical scenarios. Unfortunately, real-world graph structures inevitably contain noisy edges due to data collection errors [22], outdated information [23] or inherent variability in dynamic environments [24, 25]. These unexpected noises significantly impacts the effectiveness of GC in both the condensation and inference stages. As depicted in Fig. 1 (a), given the large training graph, the objective of the GC in condensation stage is indiscriminately emulating the distribution of the training graph, disregarding the intrinsic quality of the training graph. As a result, the condensed graph inherits the noisy distribution from the training graph, which is carried over to GNNs and compromises the model prediction accuracy [26, 27]. In addition, the representation of inductive nodes in the inference stage is more challenging with structural noises. As shown in Fig. 1 (b), inductive nodes for prediction are integrated into the large training graph only during the test time. Consequently, achieving high-quality representations of these inductive nodes becomes imperative for ensuring accurate predictions. However, the presence of noisy structures within both the training and inductive graphs propagates noisy information to these inductive nodes, exacerbating the challenge of precise representation for them. To assess the effect of noisy structures on GC, we simulate the deployment of GC under noisy training and test graphs separately. As depicted in Fig. LABEL:fig_obs (a), the performance decline on the clean test data indicates that the noisy structures in the training graph severely disrupt the graph distributions and are indiscriminately simulated by condensed graphs. Moreover, when GNNs are trained on a clean condensed graph but evaluated on the noisy inductive test graphs, the performance decline is more pronounced as shown in Fig. LABEL:fig_obs (b), underscoring the detrimental effect of noisy structures on inductive node representation. In a nutshell, the critical problem that arises for GC is: “How can GC methods effectively distill essential knowledge from noisy training graph into a condensed graph, while ensuring that GNNs trained on these condensed graphs maintain robust against noisy inductive test graphs?
@@
To empower GC with robustness against both noisy training and test graph structure, we aim to introduce a general graph structure optimization procedure in both the condensation and inference stages. Specifically, the desiderata of robust GC under the inductive setting is: (1) during condensation, the condensed graph extracted from the noisy training graph is noise-free, thus containing only the core, causal information for effective GNN training; (2) during inference, an efficient mechanism is in place to denoise the inductive graph that is only seen during test time, so as to prepare a clean graph for GNN inference on-the-fly. To do so, the first challenge comes from jointly optimizing the denoising procedure during the condensation stage. A potential solution for this challenge is graph structure learning (GSL) [28, 27], employing a specific GSL model to optimize the graph structural and learn denoised node representations. However, these GSL methods tightly entangle model training with graph structure optimization, thereby inevitably complicating GC optimization by introducing the additional model training task. Therefore, it is necessary to design an efficient and applicable optimization strategy for the denoising process. In the meantime, the second challenge involves the efficient inductive graph denoising during the inference stage. Existing GSL approaches optimize the graph structure alongside the GNN model, which restricts their capability to handle nodes observed solely during the training stage. Consequently, when inductive nodes are introduced to the training graph, these methods require retraining the model and optimize the graph from scratch, contradicting the purpose of employing GC for efficiency. Ideally, the solution would involve learning a principled denoising function during the condensation stage, capable of denoising the inductive test graph without the need for retraining.
@@
  • Through extensive experimentation, we verify that RobGC is compatible with different GC methods. It excels in handling various types and levels of graph structure noise, all while preserving the inherent generalization capabilities of GC.
@@ II-B Graph Condensation
In addition to gradient matching, a variety of GC methods adopt distinct optimization strategies to address the optimization problem in Eq. (3). For instance, GCDM [31] introduces the distribution matching, aiming to align the feature distributions between the condensed and training data. GCEM [32] directly generates the condensed graph’s eigenbasis to match that of the training graph. Meanwhile, SFGC [33] brings trajectory matching into the GC field, synchronizing the long-term learning behaviors of GNNs across the training and condensed graphs. Despite their potential to enhance GNN performance, these GC methods neglect the critical aspect of graph quality, which is essential for real-world applications.
Figure 3: The graph condensation and test-time denoising procedure of RobGC. During the training stage, the noisy training graph is condensed, and the condensed graph is utilized as a denoising signal to optimize the structure of the training graph. The condensation and denoising procedures are executed iteratively to promote mutual enhancement. In the inference stage, the noisy test graph is initially denoised by referring the condensed graph. Subsequently, well-trained GNNs are applied to this optimized test graph for inference. Dashed lines indicate noisy edges.
@@ III-B Optimization Objective and Alternating Optimization
In light of these challenges, we employ the label propagation mechanism [36] to guide the optimization procedure and determine the optimal thresholds by grid search. This training-free approach notably simplifies the optimization by eliminating hyper-parameter tuning and back-propagation procedures. To achieve this, 50% of the training nodes are randomly selected as support nodes, and their labels are propagated to predict the labels of the remaining nodes. Specifically, we construct the propagation label matrix $\bar{\mathbf{Y}}$ by retaining the one-hot labels of support nodes and marking the remaining labels as $\mathbf{0}$. Then, ${\bf A}$ is optimized based on a selected threshold candidate and $\bar{\mathbf{Y}}$ is propagated through the optimized graph structure ${\mathbb{A}}$ as:
@@ IV Experiments
We design comprehensive experiments to assess the effectiveness of RobGC and explore the following research questions: Q1: When combined with different GC methods, can RobGC maintain robust performance under varying levels of noise? Q2: Can RobGC effectively defend against various types of noise? Q3: How does RobGC perform under different condensation ratios? Q4: Does the condensed graph produced by RobGC exhibit good generalization across different GNN architectures? Q5: How does the computational efficiency of RobGC? Q6: How do the distinct components within RobGC influence its overall performance? Q7: How does the optimized graph differ from the noisy training graph? Q8: How do the different hyper-parameters affect the RobGC?
@@ IV-A Experimental Settings
Datasets. We assess the performance of RobGC on four real-world datasets: Cora, Citeseer, Pubmed [29], and Ogbn-arxiv [38]. For Cora, Citeseer, and Pubmed, we adhere to the publicly available splits for the validation and test sets, with the remaining nodes constituting the training set. Detailed information is summarized in Table I.
TABLE I: The properties of datasets. The training set is condensed in the training procedure.
Dataset #Nodes #Edges #Feature #Class Training/Validation/Test
Cora 2,708 5,429 1,433 7 1,208 / 500/ 1,000
Citeseer 3,327 4,732 3,703 6 1,827 / 500/ 1,000
Pubmed 19,717 44,338 500 3 18,217 / 500/ 1,000
Ogbn-arxiv 169,343 1,166,243 128 40 90,941/29,799/48,603

### v2

@@ Abstract
The increasing prevalence of large-scale graphs presents a significant challenge for graph neural networks (GNNs) training due to their computational demands, limiting the applicability of GNNs in various scenarios. In response to this challenge, graph condensation (GC) is proposed as a promising acceleration solution, focusing on generating an informative compact graph that enables efficient training of GNNs while retaining performance. Despite the potential to accelerate GNN training, existing GC methods overlook the quality of large training graphs during both the training and inference stages. They indiscriminately emulate the training graph distributions, making the condensed graphs susceptible to noises within the training graph and significantly impeding the application of GC in intricate real-world scenarios. To address this issue, we propose robust graph condensation (RobGC), a plug-and-play approach for GC to extend the robustness and applicability of condensed graphs in noisy graph structure environments. Specifically, RobGC leverages the condensed graph as a feedback signal to guide the denoising process on the original training graph. A label propagation-based alternating optimization strategy is in place for the condensation and denoising processes, contributing to the mutual purification of the condensed graph and training graph. Additionally, as a GC method designed for inductive graph inference, RobGC facilitates test-time graph denoising by leveraging the noise-free condensed graph to calibrate the structure of the test graph. Extensive experiments show that RobGC is compatible with various GC methods, significantly boosting their robustness.
Index Terms:
Graph condensation, graph structure learning
@@ I Introduction
Graph data [1, 2, 3] is extensively utilized across diverse domains, owing to its capability to represent complex structural relationships among various entities in the real world. To address the imperative for data management, graph neural networks (GNNs) have attracted significant attention for their exceptional representation capabilities. By incorporating node attributes and graph topology, they could extract latent information within graphs and have been utilized in a wide range of real-world applications, including chemical molecules [4], social networks [5], and recommender systems [6]. However, the burgeoning volume of graph data within these real-world applications poses formidable challenges in training GNN models. GNNs commonly employ the message-passing mechanism that propagates node features across the graph structure. Despite its effectiveness, this method can lead to the neighbor explosion issue [7], which considerably escalates the computational requirements during the model training stage. This situation becomes even more challenging when training multiple GNN models, such as in the application of neural architecture search [8], continual learning [9], and federated learning [10].
In response to the urgent demand for processing large-scale graph data, graph condensation (GC) [11, 12] is proposed to substitute the large training graph with a compact (e.g., 1,000×1,000$\times$1 , 000 × smaller) yet informative graph to accelerate the GNN training. By retaining the crucial characteristics of the large training graph, GNN models trained on the condensed graph demonstrate comparable performance to those trained on the large training graph, while significantly reducing model training time. This not only facilitates large-scale graph management but also expands the application of GNNs in resource-constrained scenarios for various downstream tasks. Among these, we specifically focus in this paper on the inductive inference task, where new and previously unseen nodes are continuously integrated into the existing training graph as depicted in Fig. 1. This task is practical and widely adopted in high-throughput graph systems [13, 14]. For instance, online social networks [15] frequently integrate new posts and users into their existing graphs for trend monitoring and anomaly detection. Large-scale e-commerce platforms [16] incorporate new products and users into their recommender systems to improve cold-start recommendations and adapt to evolving product landscapes. By integrating with GC, GNNs trained on condensed graphs can be deployed on these expanded graphs to generate embeddings for unseen nodes and make predictions [17].
Figure 2: The test accuracy of GCN [18] trained on the condensed graphs. GCond [11] is adopted to condense the graph and random noise is applied. (a) The performance of the GCN trained on condensed graphs derived from the training graph with different levels of noise and tested on a clean graph. (b) The results of the GCN trained on a clean condensed graph and tested on graphs containing varying levels of noise.
@@
Despite the potential of GC methods to expedite model training, their practical applicability remains sub-optimal in intricate real-world scenarios. Predominantly, existing GC methods presuppose the presence of clean training graph structures, contradicting the complicated nature of data processing in practical scenarios. Unfortunately, real-world graph structures inevitably contain noisy edges due to data collection errors [19], outdated information [20] or inherent variability in dynamic environments [21]. These unexpected noises significantly impacts the effectiveness of GC in both the condensation and inference stages. As depicted in Fig. 1 (a), given the large training graph, the objective of the GC in condensation stage is indiscriminately emulating the distribution of the training graph, disregarding the intrinsic quality of the training graph. As a result, the condensed graph inherits the noisy distribution from the training graph, which is carried over to GNNs and compromises the model prediction accuracy [22]. In addition, the representation of inductive nodes in the inference stage is more challenging with structural noises. As shown in Fig. 1 (b), inductive nodes for prediction are integrated into the large training graph only during the test time. Consequently, achieving high-quality representations of these inductive nodes becomes imperative for ensuring accurate predictions. However, the presence of noisy structures within both the training and inductive graphs propagates noisy information to these inductive nodes, exacerbating the challenge of precise representation for them. To assess the effect of noisy structures on GC, we simulate the deployment of GC under noisy training and test graphs separately. As depicted in Fig. 2 (a), the performance decline on the clean test data indicates that the noisy structures in the training graph severely disrupt the graph distributions and are indiscriminately simulated by condensed graphs. Moreover, when GNNs are trained on a clean condensed graph but evaluated on the noisy inductive test graphs, the performance decline is more pronounced as shown in Fig. 2 (b), underscoring the detrimental effect of noisy structures on inductive node representation. In a nutshell, the critical problem that arises for GC is: “How can GC methods effectively distill essential knowledge from noisy training graph into a condensed graph, while ensuring that GNNs trained on these condensed graphs maintain robust against noisy inductive test graphs?
@@
To empower GC with robustness against both noisy training and test graph structure, we aim to introduce a general graph structure optimization procedure in both the condensation and inference stages. Specifically, the desiderata of robust GC under the inductive setting is: (1) during condensation, the condensed graph extracted from the noisy training graph is noise-free, thus containing only the core, causal information for effective GNN training; (2) during inference, an efficient mechanism is in place to denoise the inductive graph that is only seen during test time, so as to prepare a clean graph for GNN inference on-the-fly. To do so, the first challenge comes from jointly optimizing the denoising procedure during the condensation stage. A potential solution for this challenge is graph structure learning (GSL) [23], employing a specific GSL model to optimize the graph structural and learn denoised node representations. However, these GSL methods tightly entangle model training with graph structure optimization, thereby inevitably complicating GC optimization by introducing the additional model training task. Therefore, it is necessary to design an efficient and applicable optimization strategy for the denoising process. In the meantime, the second challenge involves the efficient inductive graph denoising during the inference stage. Existing GSL approaches optimize the graph structure alongside the GNN model, which restricts their capability to handle nodes observed solely during the training stage. Consequently, when inductive nodes are introduced to the training graph, these methods require retraining the model and optimize the graph from scratch, contradicting the purpose of employing GC for efficiency. Ideally, the solution would involve learning a principled denoising function during the condensation stage, capable of denoising the inductive test graph without the need for retraining. Besides GSL, another related area is robust GNNs [24], which focus on developing specialized modules to enhance the robustness of models against noise [25] and adversarial attacks [22]. However, these model-centric approaches are constrained by the specific model architecture and downstream task, and fall short of addressing the deficiencies inherent in the data itself.
@@
  • Through extensive experimentation, we verify that RobGC is compatible with different GC methods. It excels in handling various types and levels of graph structure noise, all while preserving the inherent generalization capabilities of GC. The code is available at: https://github.com/XYGaoG/RobGC.
@@ II-B Graph Condensation
In addition to gradient matching, a variety of GC methods adopt distinct optimization strategies to address the optimization problem in Eq. (3). For instance, GCDM [27] and CAT[28] introduces the distribution matching, aiming to align the feature distributions between the condensed and training data. GDEM [29] directly generates the condensed graph’s eigenbasis to match that of the training graph. GC-SNTK [30] and OpenGC [31] leverages the kernel ridge regression to address the bi-level optimization by the close-form solution. Meanwhile, SFGC [32] and GEOM [33] brings trajectory matching into the GC field, synchronizing the long-term learning behaviors of GNNs across the training and condensed graphs. Despite their potential to enhance GNN performance, these GC methods neglect the critical aspect of graph quality, which is essential for real-world applications.
Figure 3: The training and test procedures of RobGC. During the training stage, the noisy training graph is condensed, and the condensed graph is utilized as a denoising signal to optimize the structure of the training graph. In the inference stage, the noisy test graph is denoised by referring the condensed graph. Features are omitted in test-time denoising for simplicity.
@@ III-A Reliable Graph Structure Modeling
Notably, the dual-stage structure modification does not require gradient calculations and can be performed on CPUs, enabling flexible management by processing edges and nodes in separate batches, respectively. Furthermore, this module is adaptable to incorporate more advanced sampling strategies to enhance scalability, such as prioritizing low-degree nodes that are more vulnerable in the message-passing mechanism [25]. Detailed complexity analysis is presented in Section D.
@@ III-B Optimization Objective and Alternating Optimization
In light of these challenges, we employ the label propagation mechanism [35] to guide the optimization procedure and determine the optimal thresholds by grid search. This training-free approach notably simplifies the optimization by eliminating hyper-parameter tuning and back-propagation procedures. To achieve this, 50% of the training nodes are randomly selected as support nodes, and their labels are propagated to predict the labels of the remaining nodes. [^1] We assume a clean label set to investigate structural noise. For scenarios with label noise, our method is compatible with various off-the-shelf label enhancement methods to pre-process the labels, such as ensemble learning [36], self-supervised learning [37], and large language models [38]. Specifically, we construct the propagation label matrix $\bar{\mathbf{Y}}$ by retaining the one-hot labels of support nodes and marking the remaining labels as $\mathbf{0}$. Then, ${\bf A}$ is optimized based on a selected threshold candidate and $\bar{\mathbf{Y}}$ is propagated through the optimized graph structure ${\mathbb{A}}$ as:
@@ IV Experiments
We design comprehensive experiments to assess the effectiveness of RobGC and explore the following research questions: Q1: When combined with different GC methods, can RobGC maintain robust performance under varying levels of noise? Q2: Can RobGC effectively defend against various types of noise? Q3: How does RobGC perform under different condensation ratios? Q4: Does the condensed graph produced by RobGC exhibit good generalization across different GNN architectures? Q5: How does the computational efficiency of RobGC? Q6: How do the distinct components within RobGC influence its overall performance? Q7: How does the optimized graph differ from the noisy training graph? Q8: How do the different hyper-parameters affect the RobGC? Q9: Can Robust GNNs benefit from RobGC?
@@ IV-A Experimental Settings
Datasets. We evaluate the methods on 8 datasets [11, 29] under our inductive setting, as summarized in Table I.
TABLE I: The properties of datasets.
Dataset #Nodes #Edges #Train/Val/Test nodes Train/Val/Test graph size
Cora 2708 5429 1208/500/1000 1208/1708/2208
Citeseer 3327 4732 1827/500/1000 1827/2327/2827
Pubmed 19717 44338 18217/500/1000 18217/18717/19217
Ogbn-arxiv 169343 1166243 90941/29799/48603 90941/120740/139544
Flickr 89250 899756 44625/22312/22313 44625/66937/66938
Reddit 232965 57307946 153932/23699/55334 153932/177631/209266
Squirrel 5201 396846 3701/500/1000 3701/4201/4701
Gamers 168114 13595114 84056/42028/42030 84056/126084/126086